OpenStudy (anonymous):

The ramp shown below is used to move crates of apples to loading docks at different heights. When the horizontal distance AB is 12 feet, the height of the loading dock,BC, is 6 feet. What is the height of the loading dock

OpenStudy (anonymous):

here is pic

OpenStudy (unanimoose):

I assume you mean the height of ED so: They are the exact same shape, so they have the exact same ratio. You find the ratio of 22 feet and 12 feet and you can use that to find ED. So: \[\frac{ 22 }{ 12 } \approx 1.8\]

OpenStudy (anonymous):

thx

OpenStudy (unanimoose):

Now multiply that by 6 and you will have the height, but since it is an infinite decimal it is only the >approximate< height, not exactly.

OpenStudy (unanimoose):

6 * 1.8 = height of ED.

OpenStudy (anonymous):

10.8

OpenStudy (unanimoose):

Yup!

OpenStudy (anonymous):

thx

OpenStudy (unanimoose):

Np.

OpenStudy (anonymous):

you think you could help with one more

OpenStudy (unanimoose):

Sure.

OpenStudy (anonymous):

A bag contains 10 white golf balls and 6 striped golf balls. A golfer wants to add 112 golf balls to the bag. He wants the ratio of white to striped golf balls to remain the same. How many of each should he add?

OpenStudy (unanimoose):

There are two different types, and they are going to have the same ratio so we will be adding half of 112 as white and half as striped. \[\frac{ 112 }{ 2 }\] = answer

OpenStudy (anonymous):

56?

OpenStudy (unanimoose):

Yes.

OpenStudy (anonymous):

awesome thx for the help bye.

OpenStudy (unanimoose):

Np.